PITMinutes 01-30-2017

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, K. Baggett, A. Freyberger, R. Geng, C. Hovater, A. Hutton, G. Krafft, A. Kimber, F. Marhauser, C. Mounts, R. Nelson, T. Plawski, and T. Reilly

Agenda and Notes

1. Operations Status and Summary

Physics running stabilized again. There is not much gradient margin available, especially in the NLINAC. Trip rate OK, but LEM indicates 5/hr, actually observing 8 or 9/hr.

2. SRF Hot and Cold Topics (Reece)

Feb. 6 Same "Bat-Time", Same "Bat-location"

3. Tactical PIT (Ken Baggett)

Status Presentation

Arne's Energy Reach Summary Document

Problems in the newly installed cryomodule. Stimulated tuner discussion. Bad matching between drive and stepper motors. Meeting Friday scheduled to try to get closure.

4. C57 (Reilly and Marhauser)

Two cavities meet the spec and will be put in new cryomodule. Series 001 and 003 look OK, 14.5 MV. Specs in Pansophy. Current plan C57 goes into weakest slot, probably in NLINAC. Qext came out a little off but tuned to nominal 2-3 e7.

Make sure the two "57" cavities have the best klystrons possible.

5. Regular reporting; Josh, Tom, and Thomas in two weeks.

Girder Particulate Summary

Action Items